OpenStudy (anonymous):

find the intervals on which the graph is concaving upward and downward and the inflection point for f(x)=In(x^2+6x+13

OpenStudy (kira_yamato):

Can you use the equation function, please... Arigatou gozaimasu...

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

is it \[\ln(x^2+6x+13)\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(-\infty , -5)\] U \[(1,\infty)\] : concave upward \[(-5,-1)\] : concave downward

OpenStudy (anonymous):

points of inflection would be -5, -1

OpenStudy (anonymous):

how did you get the -5 and -1 do you mind showing me mathematically

OpenStudy (anonymous):

differentiate the given function twice and equate it to zero, you'll get it after differentiating twice, we get \[(-2x^2 -12x-10)/(x^2+6x+13)^2\] equate the numerator to 0 and you should get there